OpenStudy (anonymous):

Determine how much money will be in a savings account if the initial deposit was $750 and the interest rate is 4% compounded continuously for 15 years, 6 months.

OpenStudy (anonymous):

Do you know which formula to apply?

OpenStudy (anonymous):

\[A = Pe ^{r ^{t}}\]

OpenStudy (anonymous):

Great :) So what is P?

OpenStudy (anonymous):

750?

OpenStudy (anonymous):

Yep :) r? t?

OpenStudy (anonymous):

r = .04 and for t I just multiplied 15(12) and added 6 to get 186

OpenStudy (anonymous):

You're pretty good :) Why don't you try to plug them into your calculator !

OpenStudy (anonymous):

Hold on, I believe go by year make more sense!

OpenStudy (anonymous):

6 month = .5 yr => t = 15.5

OpenStudy (anonymous):

Well that's where I go wrong somehow and I don't know where. I got $12,099.43. I put that in as the answer on my hw and it says i'm wrong.

OpenStudy (anonymous):

= 1,394

OpenStudy (anonymous):

I have to round to the nearest cent. But even when I refer to the example in the book and plug it into my calculator, I don't get the same answer. Is there a certain mode the calculator has to be in maybe? I understand how to do the problem, but I'm not coming up with the correct answer on the calculator.

OpenStudy (anonymous):

1, 394.196 = 1,394.20

OpenStudy (anonymous):

Round off to cent = check the 3rd digit after decimal point: If it's ≥ 5, then increase the "cent" digit up!

OpenStudy (anonymous):

Did you see how I round up the above result?

OpenStudy (anonymous):

For some reason I'm not even getting the thousandths place correct. I get 12099. Here's what I plugged into my calculator exactly: 750e^(.04)(15.5)

OpenStudy (anonymous):

If so, be cautious by plugging step by step: .04 * 15.5 = e^ (.62) = 750 * .. =

OpenStudy (anonymous):

I got it now . I was trying to do it as one whole step

OpenStudy (anonymous):

If you do the whole step, it'll need more parentheses! Howerver, I'm not guarantee you know where to add them, so just be safe!

OpenStudy (anonymous):

Okay, thank you very much. I feel better doing it step by step

OpenStudy (anonymous):

I'm glad that you feel better :)